OpenStudy (anonymous):

An object is projected at u m/s at elevation angle \(\theta\) at a height of 10m. The object hits the ground 100m away. Suppose \(\theta\) = \(tan^{-1}\left(\frac{3}{4}\right)\) Find the velocity \(u\)

OpenStudy (anonymous):

This what I did: \[Max~ height: \frac{dy}{dt} = -9.8t + \frac{3}{5} u =0 => t = \frac{3u}{49} \] \[Flight time: 2* t ~to ~reach~ the \max. height => 2*\frac{3u}{49} = \frac{6u}{49} \] \[x = \left(\frac{4}{5}u\right) t => x = \left(\frac{4}{5}u\right) * \frac{6u}{49} => \frac{24u^{2}}{245} = 100 => u = 31..\theta \]

OpenStudy (anonymous):

\(u = 31..\)

OpenStudy (anonymous):

Where did i go wrong?

OpenStudy (anonymous):

check problem flight line

OpenStudy (anonymous):

flight time?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you take the 2 times the faction

OpenStudy (anonymous):

i dont think there is anything wrong with it :/ flight time = 2 * time to reach the max. height

OpenStudy (anonymous):

2* (3u/49) => ( 6u/49)

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

u = 30

OpenStudy (anonymous):

okay so help me to understand your last line and why there are two 4/5c in this line for?

OpenStudy (anonymous):

where did that come from? I don't see it?

OpenStudy (anonymous):

the first x = (4/5)ut is the question then i subbed the t in which is equal to => x = (4/5)u*(6u/49)

OpenStudy (anonymous):

i mean equation not question lol

OpenStudy (anonymous):

but you have it twice not once

OpenStudy (anonymous):

that is my question where or how did the second 4/5c come into play?

OpenStudy (anonymous):

easier to understand now \[=> x = \left(\frac{4}{5}u\right) * \frac{6u}{49} => \frac{24u^{2}}{245} = 100 => u = 31\]

OpenStudy (anonymous):

that is the range \[=> x = \left(\frac{4}{5}u\right) t\]

OpenStudy (anonymous):

please bare with me I am trying to see where you are off is all

OpenStudy (anonymous):

ok thanks. i have no idea where i went wrong; it looks right to me

OpenStudy (anonymous):

yes you are right, although being off one is odd..

OpenStudy (anonymous):

where did i go wrong?

OpenStudy (anonymous):

that is why I am asking questions to you nothing more

OpenStudy (anonymous):

hey I think the best way know where you went wrong is to start again. May be the mistake is small and can not be seen otherwise

OpenStudy (anonymous):

well i did the several times and got the same results that is why i posted here..

OpenStudy (anonymous):

are you working from a text book?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

where did this problem and answer come from than?

OpenStudy (anonymous):

from my notes ; i dont think that its a problem from a textbook

OpenStudy (anonymous):

how do know than it's 30 not 31 than?? help to understand I am lost here if not from a text book and only from your notes? just trying to understand this better

OpenStudy (anonymous):

idea check your notes again and make sure you copied the answer right. It could be that you are right and the answer you copied is just wrong. just a thought

OpenStudy (anonymous):

well; its a problem from my teachers set of notes that he gave to me; which i believe he made it himself so its not from a textbook. And the answer is right it cannot be wrong.

OpenStudy (anonymous):

forget it about it. thanks anyway

OpenStudy (anonymous):

there is a chance the teacher may have made a mistake and didn't catch it. I have had it happen to me

OpenStudy (anonymous):

well the problem is that there is a solution in it which gives the right answer; but there are several ways of solving the problem. that is why i wanted to try another approach; which indeed failed; but i did not know why :/

OpenStudy (anonymous):

wait a second how about we take the answer and work backwards now

OpenStudy (anonymous):

to see if the answer gives us the start of the problem

OpenStudy (anonymous):

i dont need to work backwards. i just want to know why my approach went wrong

OpenStudy (anonymous):

hm, forget about this problem; i will try another approach. thanks anywy!

OpenStudy (anonymous):

no problem I am srry I can not help you :(

OpenStudy (anonymous):

CAN I HELP?

OpenStudy (anonymous):

yes please!

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):